OpenStudy (anonymous):

A solutiob of thr equation cos 2 theta + sin 2 theta = -1

OpenStudy (gorv):

cos2x+sin2x=-1

OpenStudy (gorv):

multiply each side by 1/(sqrt2)

OpenStudy (gorv):

\[\frac{ 1 }{ \sqrt{2} }*\cos2x+\frac{ 1 }{ \sqrt{2} *\sin2x}=\frac{ -1 }{ \sqrt{2} }\]

OpenStudy (gorv):

\[\frac{ 1 }{ \sqrt{2} }=\cos45=\sin45\]

OpenStudy (gorv):

\[\sin45*\cos2x+\cos45*\sin2x=\sin(45+2x)=\frac{ -1 }{ \sqrt{2} }\]

OpenStudy (gorv):

\[\frac{ \pi }{ 4 }+2x=\sin^{-1}(\frac{ -1 }{ \sqrt{2} })=-\frac{ \pi }{ 4 }\]

OpenStudy (gorv):

2x==-pi/4-pi/4=-pi/2 x=-pi/4

OpenStudy (gorv):

it can have many solution this is just one